OpenStudy (toxicsugar22):

Ether, the well known anesthetic, has a density of 0.736 g/cm^3. What is the volume, in litter, of 225.0 grams of ether?

OpenStudy (toxicsugar22):

do you undertand this

OpenStudy (psymon):

Density = mass/volume. So if the density is .736g, then we would need to solve this equation: \[0.736=\frac{ 225.0 }{ V }\]

OpenStudy (toxicsugar22):

165.6 is that it

OpenStudy (psymon):

No, it isnt actually. You need to get V out of the denominator of that fraction in order to solve for it. To get V out of the bottom, you multiply both sides by V. \[0.736V = \frac{ 225.0V }{ V }\implies 0.736V = 225.0\] Now that we have 0.736V = 225.0, we can divide both sides by 0.736 to isolate V: \[\frac{ 0.736V }{ 0.736 }=\frac{ 225.0 }{ 0.736 }\implies V = \frac{ 225.0 }{ 0.736 }\]So that would be the division you need to do.

OpenStudy (toxicsugar22):

305.7065

OpenStudy (toxicsugar22):

right

OpenStudy (psymon):

Right :3

OpenStudy (toxicsugar22):

and the units would be cm^3

OpenStudy (toxicsugar22):

right

OpenStudy (psymon):

Yes it would be :3

OpenStudy (toxicsugar22):

thanks

OpenStudy (psymon):

Mhm, np.

OpenStudy (toxicsugar22):

can you help me with one mor e

OpenStudy (psymon):

If I know how to do it then sure, lol.

OpenStudy (toxicsugar22):

an automobile manufacture claims that is sedan uses only 6.0 L of gasoline to travel 100 km. How many miles per gallon of gasoline could be expected from this car ( 1 mile = 1.61 km)

OpenStudy (psymon):

Well, this requires us to have a liters to gallons conversion, too.

OpenStudy (psymon):

It wasnt given to you?

OpenStudy (psymon):

Professor Google it is.

OpenStudy (psymon):

Okay, so 1 gallon = 3.79L

OpenStudy (toxicsugar22):

okay

OpenStudy (psymon):

So what we do is we find conversion factors, for example 1 gallon = 3.79L, 1 mile = 1.61km, etc, and we multiply those conversion factors by our original information. So what we're initially told is 6.0L can get you 100km. So what I want to do is turn this into a fraction: \[\frac{ 6.0L }{ 100km }\]Now I can decide to change km to miles or liters to gallons first, it doesnt matter. So Ill do liters to gallons first. The conversion factor is 1g = 3.79L. So now I want to turn this into a fraction and multiply it by the current fraction we have: \[\frac{ 6.0L }{ 100km }*\frac{ 1gal }{ 3.79L }\]Now the reason I wrote it the way I did is because when I multiply, I want the units to cancel. I dont want liters anymore, I want gallons. So when I multiply by the conversion factor of 1gal = 3.79L, I put the fraction as such so that Liters will cancel out top and bottom. SO doing that multiplication, that gives me: \[\frac{ 6.0gal }{ 379km }\]Now I finally want to change km to miles. So now Ill take my other conversion factor of 1 mile = 1.61km, and multiply this by what we currently have. Now, we want the km to go away, so that means when we multiply, I want to have km on top and km also on bottom. That means Im multiplying liek this: \[\frac{ 6.0gal }{ 379km}*\frac{ 1.61km }{ 1mile }\]When we multiply these, the km will cancel top and bottom and then it's just calculator work. \[\frac{ 9.66gal }{ 379mil }\]Now this gives us the amount for 9.66 gallons, but we want the amount for 1 gallon. So that requires us to set up a proportion. The proportion would be set up like this: \[\frac{ 9.66gal }{ 379mil }= \frac{ 1gal }{ x }\]So basically this says that, assumign the ratio is the same, if 9.66 gallons = 379 miles, then 1gal = x miles. This can then be solved by cross-multiplying. So 9.66 multiplies across by x and 1 gallon multiplies straight across by 379 miles. Doing the cross-multiplication gives us: \[9.66x = 379\]So then solving for x gives me: \[x=\frac{ 379 }{ 9.66 }\approx 39.23\]So that means I would get approximately 39.23 miles per gallon.

OpenStudy (toxicsugar22):

ohh wow thanks

OpenStudy (psymon):

As long as it makes sense xD

OpenStudy (toxicsugar22):

thanks again it made sense

OpenStudy (psymon):

Sure ^_^